QUESTION 9

  1. Based on an examination of a sample of 895 cases drawn from the General Social Survey, you find that the mean number of hours worked last week was 41.47, with a standard deviation of 15.04.

    a. What is the probability that someone in the sample worked 60 hours or more?
    b. How many people worked 60 hours or more?
Answer #1

a)
P(X >= 60) = P(z <= (60 - 41.47)/15.04)
= P(z >= 1.23)
= 1 - 0.8907
= 0.1093

b)
number of people = 895 * 0.1093 = 97.82
If rounded to interger value, no. of people are 98